# Online Conformal Prediction with Efficiency Guarantees

### v1

@@ 1 Introduction
Achieving (1) is referred to as coverage. In addition to coverage, we would like a conformal predictor to be efficient, in that it is not too large. For example, a trivial predictor that always outputs the whole space $C=\mathcal{Y}$, will achieve coverage, but not necessarily be efficient. The most commonly studied settings by far are where $\mathcal{Y}$ is either a discrete label space, or $\mathcal{Y}=[0,1]$. [^1] Most conformal methods are designed via a black-box score function which assigns a (potentially high-dimensonial) datapoint a real-valued score, and operating on the quantiles of the scores of the datapoints. The conformal procedures are then designed over the quantile space $[0,1]$. We will focus on the setting where $\mathcal{Y}=[0,1]$. Conformal prediction is a well-studied problem in statistics, and we refer the reader to the book of Angelopoulos et al. (2025) for a full introduction.
The assumption that the test point is chosen uniformly at random is called exchangeability. This assumption is motivated by the goal of conformal prediction is to produce statistically valid predictions under the fewest possible statistical assumptions. A typical application of conformal prediction is in uncertainty quantification. For example, suppose the $Y_{i}$ are the errors of a black-box machine learning model on different data points. Then we can think of the size of a conformal set $C$ for $\mathbf{Y}$ as a measure of the uncertainty of the model. In this setting, we would like to avoid making assumptions about the distribution of model error and data. For example, we may not be able to assume that the data are independent.
@@
It is clear that the type of “one step” guarantee that is typically studied in conformal prediction is impossible. If the $Y_{i}$s are fully arbitrary, then a conformal predictor essentially has to output the whole space $\mathcal{Y}$ [^2] More formally, the predictor has to cover at least a $(1-\alpha)$-fraction of $[0,1]$ in expectation, which is achieved by a trivial predictor which ignores the training data and outputs a random subset of $[0,1]$. However, in an online setting, we can aim to instead achieve coverage on average over time.
@@
and our goal is to design an algorithm with low volume approximation (competitive ratio) $\mu$, where
@@ 1.1 Results
for all permutations $\sigma:[T]\rightarrow[T]$, where =$\overset{\mathrm{d}}{=}$ means they are equal in distribution. This is also referred to as the random order model. Note that the class of exchangeable sequences generalizes the class of i.i.d. sequences, as i.i.d. sequences are always exchangeable, but an exchangeable sequence need not be i.i.d..
@@
Is it possible to design an algorithm for fully arbitrary data that recovers the best known bounds for exchangeable data?
@@
3:     if $I_{\mathrm{current}}$ has empirical coverage $<1-R(t-1)$ over $\{y_{t^{\prime}}:t^{\prime}<t\}$ then
@@
[…]
@@ 1.2 Technical Overview
One of the beautiful properties of classical online learning is that it is possible to design algorithms for arbitrary input sequences that have vanishing regret, i.e., they make $1+O(\sqrt{\log n/T})$ times as many errors as the best fixed expert in hindsight (see e.g., Blum, Hopcroft, and Kannan, 2020). This vanishing regret term matches the sampling error that one would face even for i.i.d. sequences. In essence, this says that going from i.i.d. (and therefore exchangeable) data to arbitrary data is without loss, as we are able to achieve the best possible guarantee even in the arbitrary case. Hence there is one algorithm that is able to achieve the “best of both worlds.” Such results are also known in bandit problems and other settings (see e.g., Bubeck and Slivkins, 2012).
@@
Designing an algorithm for arbitrary input sequences presents some challenges. Consider the following example. Our goal is to design a conformal predictor for miscoverage rate $\alpha=\frac{1}{2}$. There are two sequences. Let $\varepsilon>0$ be some small parameter. $S^{(1)}$ is drawn uniformly from $[0,\varepsilon]$ for the first half of the sequence, and then expands in scale to be drawn uniformly from $[0,1]$ for the second half of the sequence. $S^{(2)}$ is also drawn uniformly from $[0,\varepsilon]$ for the first half of the sequence, but then shrinks in scale to be drawn uniformly from $[0,\varepsilon^{2}]$ for the second half of the sequence. That is,
$S^{(1)}=(S^{(1)}_{1},\dots,S^{(1)}_{T}),\quad$ $S^{(1)}_{t}\sim\begin{cases}\mathrm{Unif}[0,\varepsilon]&\text{for }t\leq T/2\\ \mathrm{Unif}[0,1]&\text{for }t>T/2\end{cases},$
$S^{(2)}=(S^{(2)}_{1},\dots,S^{(2)}_{T}),\quad$ $S^{(2)}_{t}\sim\begin{cases}\mathrm{Unif}[0,\varepsilon]&\text{for }t\leq T/2\\ \mathrm{Unif}[0,\varepsilon^{2}]&\text{for }t>T/2\end{cases}.$
@@
However, there is a more significant issue. Even if $\mathsf{minwidth}$ is fixed, the above example still forces any algorithm that captures a constant fraction of points to incur a multiplicative error $\Omega(1/\mathsf{minwidth})$ in volume, which one should think of as being extremely large. This issue arises because we require the algorithm to achieve coverage very close to the target, i.e., a non-trivial guarantee for $\alpha=\frac{1}{2}$ requires the algorithm to make $<T=2\alpha T$ mistakes. Surprisingly, the problem actually becomes tractable for smaller values of $\alpha$ if we consider a biciteria approximation, that competes with the best interval that makes at most $\alpha T$ mistakes, while making a multiplicative factor more than $\alpha T$ mistakes.
@@
The analysis of the algorithm for exchangeable sequences is tricky because the interval $I_{t}$ that Algorithm 1 plays on day $t$ is dependent on the first $t-1$ days of the sequence $S_{1},\dots,S_{t-1}$. This is in turn not independent of $S_{t}$. Thus the coverage on day $t$
@@
To understand the challenge, consider the following example. Let $S$ be a uniformly random permutation of a multiset that has $(1-2\alpha)T$ copies of $1/2$, $\alpha T$ copies of 00 and $\alpha T$ copies of $1$. $S$ is exchangeable by construction. On day $T$, we have seen all but one element, which could have been $1/2$ with probability $1-2\alpha$, 00 with probability $\alpha$, or $1$ with probability $\alpha$.
Algorithm 1, for $\mu=1$, $R(t)=\alpha-O(\sqrt{\log T/T})$, will arbitrarily choose a minimum volume interval that achieves coverage $1-R(T-1)$ on the first $T-1$ points. The intervals $[0,1/2]$ or $[1/2,1]$ will both meet the coverage requirement and have equal (minimal) volume, so say that Algorithm 1 tiebreaks by choosing the interval that achieved higher coverage, and choosing $[0,1/2]$ if they have equal coverage. Then, if $S_{T}=0$, $[1/2,1]$ will achieve higher coverage on the prefix and $S_{T}$ will not be in the output interval. Similarly, if $S_{T}=1$, $[0,1/2]$ will achieve higher coverage on the prefix and $S_{T}$ will not be in the output interval. Thus $S_{T}$ is covered with probability $\leq 1-2\alpha$, when we are aiming for coverage $1-\alpha$. Our analysis handles this subtle issue by showing that there cannot be many days $t$ on which this issue arises. This allows us to amortize the loss in coverage by allowing for higher error rates both at the beginning and at the end of the sequence. Carefully accounting for this allows us to show that the standard analysis for i.i.d. sequences goes through even for sequences that are only exchangeable and not necessarily i.i.d..
@@ 2 Algorithmic Guarantees for Arbitrary Sequences
Now we bound the number of mistakes. Let $t_{1},t_{2}$ be two subsequent times that $I_{\mathrm{current}}$ is reset. That is, the same $I_{\mathrm{current}}$ is played on all days $[t_{1},t_{2})$. Call $[t_{1},t_{2})$ a phase. We begin by bounding the number of mistakes in a phase. Since $I_{\mathrm{current}}$ is feasible on day $t_{2}-1$, $I_{\mathrm{current}}$ makes $\leq R(t_{2}-2)(t_{2}-2)$ mistakes on $\{y_{1},\dots,y_{(t_{2}-1)}\}$. Therefore $I_{\mathrm{current}}$ makes $\leq R(t_{2}-2)(t_{2}-2)+1$ mistakes on this phase: $\{y_{t_{1}},\dots,y_{t_{2}-1}\}$. Thus the number of mistakes that the algorithm makes on any phase is
@@
Partition the sequence after day $t^{\star}$ into $O(\log T)$ epochs, where each epoch $[t_{\mathrm{start}},t_{\mathrm{end}}]$ has $\frac{t_{\mathrm{end}}-1}{t_{\mathrm{start}}-1}<3$. We bound the number of mistakes that Algorithm 1 makes on each epoch. Fix an epoch $e$, and consider the first time in this epoch that Algorithm 1 resets $I_{\mathrm{current}}$. By line 9, $I_{\mathrm{current}}$ is set to have volume $\geq\mu\mathsf{minwidth}$.
@@
$=t^{\star}+O\left(\frac{\log(1/\mathsf{minwidth})\log T}{\log(\mu)}\left(1+\max_{0\leq t\leq T-1}R(t)t\right)\right).$
@@
For the bound on number of mistakes, this choice of $R(t)$ ensures that an interval is feasible on day $t$ if and only if it has made $\leq\alpha T$ mistakes up to day $t$. Consider $t_{1},t_{2}$ such that $2\alpha T+1<t_{1}<t_{2}$. Let $I_{1}$ be any feasible interval on day $t_{1}$ and $I_{2}$ be any feasible interval on day $t_{2}$. Since $I_{1}$ and $I_{2}$ each make $\leq\alpha T$ mistakes over the first $12\alpha T+$ 1 days, they must capture at least one day in common. Therefore $I_{1}$ and $I_{2}$ overlap. Thus we can think of all days $t>2\alpha T+1$ as forming one epoch, and the proof of Theorem 2.1 (Equation 4) tells us that the number of phases in this epoch is
$=O\left(\frac{\log(1/\mathsf{minwidth})}{\log(\mu)}\right).$
@@ 3 Algorithmic Guarantees for Exchangeable Sequences
Quantitatively, we can access this property through uniform convergence. In the most commonly used form, uniform convergence tells us that for a set family $\mathcal{F}$ of bounded VC-dimension, for us the family of intervals over $[0,1]$, and a set $Y$ of $n$ samples drawn i.i.d. from some distribution $\mathcal{D}$, the coverage of every set in $\mathcal{F}$ over $Y$ simultaneously converges to its true coverage over $\mathcal{D}$. That is, treating the VC-dimension of $\mathcal{F}$ and failure probability as constants, the coverage of every set in $\mathcal{F}$ over $Y$ is within ±$\pm O(1/\sqrt{n})$ of its coverage over $\mathcal{D}$.
@@
Uniform convergence for exchangeable sequences tells us that, treating the VC-dimension of $\mathcal{F}$ and failure probability as constants, the coverage of every set in $\mathcal{F}$ over $Y$ is within ±$\pm O(1/\sqrt{t})$ of its coverage over $\mathbf{S}$. That is, the coverage of sets in $\mathcal{F}$ over subsequences of $\mathbf{S}$ behaves essentially as though the elements of $Y$ were drawn i.i.d. from $\mathrm{Unif}\{\mathbf{S}_{1},\dots,\mathbf{S}_{T}\}$. A formal version of this statement can be found in Lemma 6.1.
@@
We analyze the algorithm’s performance on one particular day $t$. We will refer to the days before $t$, i.e. the window $[1,t-1]$, as the prefix, and the days $t$ and after, i.e. the window $[t,T]$, as the suffix. For an interval $I\subseteq[0,1]$ and a window (contiguous subsequence) $W=[t_{1},t_{2}]\subseteq[1,T]$, define $\mathrm{err}_{\mathbf{S}}(I,W)$ to be the fraction of days in $W$ of $\mathbf{S}$ that are not covered by $I$. Since the set of intervals over $[0,1]$ has VC-dimension 2, Lemma 6.1 tells us that the following statements hold, each with probability $\geq 1-\frac{1}{T^{3}}$, for some absolute constant $C$:
@@
Consider any interval $I$ that has coverage $\geq 1-\min_{0\leq t\leq T}\left[R(t)+C\sqrt{\log T/t}\right]$ over $\mathbf{S}$ in hindsight. Then, conditioned on (5) holding for all $t$, $I$ must be a feasible choice for $I_{t}$ on every day $t$ (Algorithm 1, Line 6). Thus, conditioned on (5) holding for all $t$, Algorithm 1 plays intervals of maximum volume
@@
Now let $\widehat{S}$ be any realization of $\mathbf{S}$ that satisfies the uniform convergence property. Let $K$ be the number of times sets $I_{\mathrm{current}}$ Algorithm 1 (for $\mu$, $\mathsf{minwidth}$, $R(t)=\frac{\alpha T}{t}$) on $\widehat{S}$. Let $t_{i}$ for $1\leq i\leq K$ be the days when $I_{\mathrm{current}}$ is set, where $t_{1}=1$, and we define $t_{K+1}=T+1$ for convenience. We refer to the window $[t_{i},t_{i-1})$ as phase $i$.
Let $M^{\widehat{S}}$ be the number of mistakes that Algorithm 1 makes on $\widehat{S}$. Let $M^{\widehat{S}}_{i}$ be the miscoverage rate of Algorithm 1 over phase $i$ of $\widehat{S}$. Let $I_{i}$ be the value of $I_{\mathrm{current}}$ on phase $i$. If Algorithm 1 resets $I_{\mathrm{current}}$ on day $t_{i+1}$, it means that $I_{i}$ has error rate $\leq\frac{\alpha T}{t_{i+1}-1}$ over the window $[1,t_{i+1}-1]$ (Algorithm 1, Line 3). By the uniform convergence property (9),

### v2

@@ 1 Introduction
Achieving (1) is referred to as coverage. In addition to coverage, we would like a conformal predictor to be efficient, in that it is not too large. For example, a trivial predictor that always outputs the whole space $C=\mathcal{Y}$, will achieve coverage, but not necessarily be efficient. The most commonly studied settings by far are where $\mathcal{Y}$ is either a discrete label space, or $\mathcal{Y}=[0,1]$. We will focus on the setting where $\mathcal{Y}=[0,1]$. Conformal prediction is a well-studied problem in statistics, and we refer the reader to the book of Angelopoulos, Barber, and Bates (2025) for a full introduction.
The assumption that the test point is chosen uniformly at random is called exchangeability. This assumption is motivated by the goal of conformal prediction is to produce statistically valid predictions under the weakest possible statistical assumptions. A typical application of conformal prediction is in uncertainty quantification. For example, suppose the $Y_{i}$ are the errors of a black-box machine learning model on different data points. Then we can think of the size of a conformal set $C$ for $\mathbf{Y}$ as a measure of the uncertainty of the model. In this setting, we would like to avoid making assumptions about the distribution of model error and data. For example, we may not be able to assume that the data are independent.
@@
It is clear that the type of “one step” guarantee that is typically studied in conformal prediction is impossible. If the $Y_{i}$s are fully arbitrary, then a conformal predictor essentially has to output the whole space $\mathcal{Y}$. [^1] More formally, the predictor has to cover at least a $(1-\alpha)$-fraction of $[0,1]$ in expectation, which is achieved by a trivial predictor which ignores the training data and outputs a random subset of $[0,1]$. However, in an online setting, we can aim to instead achieve coverage on average over time.
@@
and our goal is to design an algorithm with low volume approximation (competitive ratio) $\mu$, where we define
@@ 1.1 Results
for all permutations $\sigma:[T]\rightarrow[T]$, where $\overset{\mathrm{d}}{=}$ means they are equal in distribution. This is also referred to as the random order model. Note that the class of exchangeable sequences generalizes the class of i.i.d. sequences, as i.i.d. sequences are always exchangeable, but an exchangeable sequence need not be i.i.d..
@@
Is it possible to design an algorithm for fully arbitrary data that recovers the best possible bounds for exchangeable data?
@@
3:  if $I_{\mathrm{current}}$ has empirical coverage <1R(t1)<1-R(t-1) over $\{y_{t^{\prime}}:t^{\prime}<t\}$ then
@@
     $\text{ for any }\varepsilon^{\prime}>0,\quad\Omega\left(\min\left\{\frac{\log(1/\mathsf{minwidth})}{\log(\mu_{\mathrm{avg}})}\alpha^{1+\varepsilon^{\prime}}T,~T\right\}\right)~\text{mistakes}$
@@ 1.2 Technical Overview
One of the beautiful properties of classical online learning is that it is possible to design algorithms for arbitrary input sequences that have vanishing regret, i.e., they make $1+O(\sqrt{\log n/T})$ times as many errors as the best fixed expert in hindsight (see e.g., Blum, Hopcroft, and Kannan (2020)). This vanishing regret term matches the sampling error that one would face even for i.i.d. sequences. In essence, this says that going from i.i.d. (and therefore exchangeable) data to arbitrary data is without loss, as we are able to achieve the best possible guarantee even in the arbitrary case. Hence there is one algorithm that is able to achieve the “best of both worlds.” Such results are also known in bandit problems (Bubeck and Slivkins, 2012), online conformal prediction (without efficiency guarantees) (Angelopoulos, Barber, and Bates, 2024), and other settings.
@@
Designing an algorithm for arbitrary input sequences presents some challenges. Consider the following example. Our goal is to design a conformal predictor for miscoverage rate $\alpha=\frac{1}{2}$. There are two sequences. Let $\varepsilon>0$ be some small parameter. $S^{(1)}$ is drawn uniformly from $0,\varepsilon]$ for the first half of the sequence, and then expands in scale to be drawn uniformly from $[0,1]$ for the second half of the sequence. $S^{(2)}$ is also drawn uniformly from $0,\varepsilon]$ for the first half of the sequence, but then shrinks in scale to be drawn uniformly from $0,\varepsilon^{2}]$ for the second half of the sequence. That is,
$S^{(1)}=(S^{(1)}_{1},\dots,S^{(1)}_{T})$$\quad$ $S^{(1)}_{t}\sim\begin{cases}\mathrm{Unif}[0,\varepsilon]&\text{for }t\leq T/2\\ \mathrm{Unif}[0,1]&\text{for }t>T/2\end{cases},$
$S^{(2)}=(S^{(2)}_{1},\dots,S^{(2)}_{T})$$\quad$ $S^{(2)}_{t}\sim\begin{cases}\mathrm{Unif}[0,\varepsilon]&\text{for }t\leq T/2\\ \mathrm{Unif}[0,\varepsilon^{2}]&\text{for }t>T/2\end{cases}.$
@@
However, there is a more significant issue. Even if $\mathsf{minwidth}$ is fixed, the above example still forces any algorithm that captures a constant fraction of points to incur a multiplicative error $\Omega(1/\mathsf{minwidth})$ in volume, which one should think of as being extremely large. This issue arises because we require the algorithm to achieve coverage very close to the target, i.e., a non-trivial guarantee for $\alpha=\frac{1}{2}$ requires the algorithm to make $T=2\alpha T$ mistakes. Surprisingly, the problem actually becomes tractable for smaller values of $\alpha$ if we consider a biciteria approximation, that competes with the best interval that makes at most $\alpha T$ mistakes, while making a multiplicative factor more than $\alpha T$ mistakes.
@@
The analysis of the algorithm for exchangeable sequences is tricky because the interval $I_{t}$ that Algorithm 1 plays on day $t$ is dependent on the first t1t-1 days of the sequence $S_{1},\dots,S_{t-1}$. This is in turn not independent of $S_{t}$. Thus the coverage on day $t$
@@
To understand the challenge, consider the following example. Let $S$ be a uniformly random permutation of a multiset that has $(1-2\alpha)T$ copies of $1/2$, $\alpha T$ copies of 0 and $\alpha T$ copies of $1$. $S$ is exchangeable by construction. On day $T$, we have seen all but one element, which could have been $1/2$ with probability $1-2\alpha$, 0 with probability $\alpha$, or $1$ with probability $\alpha$.
Algorithm 1, for $\mu=1$, $R(t)=\alpha-O(\sqrt{\log T/T})$, will arbitrarily choose a minimum volume interval that achieves coverage 1R(T1)1-R(T-1) on the first T1T-1 points. The intervals $[0,1/2]$ or $[1/2,1]$ will both meet the coverage requirement and have equal (minimal) volume, so say that Algorithm 1 tiebreaks by choosing the interval that achieved higher coverage, and choosing $[0,1/2]$ if they have equal coverage. Then, if $S_{T}=0$, $[1/2,1]$ will achieve higher coverage on the prefix and $S_{T}$ will not be in the output interval. Similarly, if $S_{T}=1$, $[0,1/2]$ will achieve higher coverage on the prefix and $S_{T}$ will not be in the output interval. Thus $S_{T}$ is covered with probability $\leq 1-2\alpha$, when we are aiming for coverage $1-\alpha$. Our analysis handles this subtle issue by showing that there cannot be many days $t$ on which this issue arises. This allows us to amortize the loss in coverage by allowing for higher error rates both at the beginning and at the end of the sequence. Carefully accounting for this allows us to show that the standard analysis for i.i.d. sequences goes through even for sequences that are only exchangeable and not necessarily i.i.d..
@@ 1.3 Related Work
”Best of both worlds” for Online Conformal Prediction.
We note that the work of Angelopoulos, Barber, and Bates (2024) provides a “best of both worlds” guarantee for online conformal prediction. This does not contradict our result that there is no possible “best of both worlds” guarantee, as the two works are referring to different notions of optimality. Angelopoulos, Barber, and Bates (2024) provide a guarantee that simultaneously achieves average coverage tending to $1-\alpha$ on arbitrary sequences, and coverage on each day tending to $1-\alpha$ on exchangeable sequences. In this work, we are interested in optimizing efficiency. We consider algorithms that achieve average coverage both for arbitrary and exchangeable sequences, and we would like to show that they achieve the best possible volume approximation ratio in these two settings. We show that this type of guarantee is impossible to achieve.
@@ 2 Algorithmic Guarantees for Arbitrary Sequences
Now we bound the number of mistakes. Let $t_{1},t_{2}$ be two subsequent times that $I_{\mathrm{current}}$ is reset. That is, the same $I_{\mathrm{current}}$ is played on all days $t_{1},t_{2})$. Call $t_{1},t_{2})$ a phase. We begin by bounding the number of mistakes in a phase. Since $I_{\mathrm{current}}$ is feasible on day $t_{2}-1$, $I_{\mathrm{current}}$ makes $\leq R(t_{2}-2)(t_{2}-2)$ mistakes on $\{y_{1},\dots,y_{(t_{2}-1)}\}$. Therefore $I_{\mathrm{current}}$ makes $\leq R(t_{2}-2)(t_{2}-2)+1$ mistakes on this phase: $\{y_{t_{1}},\dots,y_{t_{2}-1}\}$. Thus the number of mistakes that the algorithm makes on any phase is
@@
Partition the sequence after day $t^{\star}$ into $O(\log T)$ epochs, where each epoch $t_{\mathrm{start}},t_{\mathrm{end}}]$ has $\frac{t_{\mathrm{end}}-1}{t_{\mathrm{start}}-1}<3$. We bound the number of mistakes that Algorithm 1 makes on each epoch. Fix an epoch $e$, and consider the first time in this epoch that Algorithm 1 resets $I_{\mathrm{current}}$. By line 9, $I_{\mathrm{current}}$ is set to have volume $\geq\mu\mathsf{minwidth}$.
@@
$t^{\star}+O\left(\frac{\log(1/\mathsf{minwidth})\log T}{\log(\mu)}\left(1+\max_{0\leq t\leq T-1}R(t)t\right)\right).$
@@
For the bound on number of mistakes, this choice of $R(t)$ ensures that an interval is feasible on day $t$ if and only if it has made $\leq\alpha T$ mistakes up to day $t$. Consider $t_{1},t_{2}$ such that $2\alpha T+1<t_{1}<t_{2}$. Let $I_{1}$ be any feasible interval on day $t_{1}$ and $I_{2}$ be any feasible interval on day $t_{2}$. Since $I_{1}$ and $I_{2}$ each make $\leq\alpha T$ mistakes over the first $2\alpha T+1$ days, they must capture at least one day in common. Therefore $I_{1}$ and $I_{2}$ overlap. Thus we can think of all days $t>2\alpha T+1$ as forming one epoch, and the proof of Theorem 2.1 (Equation 4) tells us that the number of phases in this epoch is
$O\left(\frac{\log(1/\mathsf{minwidth})}{\log(\mu)}\right).$
@@ 3 Algorithmic Guarantees for Exchangeable Sequences
Quantitatively, we can access this property through uniform convergence. In the most commonly used form, uniform convergence tells us that for a set family $\mathcal{F}$ of bounded VC-dimension, for us the family of intervals over $[0,1]$, and a set $Y$ of $n$ samples drawn i.i.d. from some distribution $\mathcal{D}$, the coverage of every set in $\mathcal{F}$ over $Y$ simultaneously converges to its true coverage over $\mathcal{D}$. That is, treating the VC-dimension of $\mathcal{F}$ and failure probability as constants, the coverage of every set in $\mathcal{F}$ over $Y$ is within $\pm O(1/\sqrt{n})$ of its coverage over $\mathcal{D}$.
@@
Uniform convergence for exchangeable sequences tells us that, treating the VC-dimension of $\mathcal{F}$ and failure probability as constants, the coverage of every set in $\mathcal{F}$ over $Y$ is within $\pm O(1/\sqrt{t})$ of its coverage over $\mathbf{S}$. That is, the coverage of sets in $\mathcal{F}$ over subsequences of $\mathbf{S}$ behaves essentially as though the elements of $Y$ were drawn i.i.d. from $\mathrm{Unif}\{\mathbf{S}_{1},\dots,\mathbf{S}_{T}\}$. A formal version of this statement can be found in Lemma 6.1.
@@
We analyze the algorithm’s performance on one particular day $t$. We will refer to the days before $t$, i.e. the window [1,t1][1,t-1], as the prefix, and the days $t$ and after, i.e. the window $[t,T]$, as the suffix. For an interval $I\subseteq[0,1]$ and a window (contiguous subsequence) $W=[t_{1},t_{2}]\subseteq[1,T]$, define $\mathrm{err}_{\mathbf{S}}(I,W)$ to be the fraction of days in $W$ of $\mathbf{S}$ that are not covered by $I$. Since the set of intervals over $[0,1]$ has VC-dimension 2, Lemma 6.1 tells us that the following statements hold, each with probability $\geq 1-\frac{1}{T^{3}}$, for some absolute constant $C$:
@@
Consider any interval $I$ that has coverage $\geq 1-\min_{0\leq t\leq T}\left[R(t)+C\sqrt{\log T/t}\right]$ over $\mathbf{S}$ in hindsight. Then, conditioned on (5) holding for all $t$, II must be a feasible choice for $I_{t}$ on every day $t$ (Algorithm 1, Line 6). Thus, conditioned on (5) holding for all $t$, Algorithm 1 plays intervals of maximum volume
@@
Now let $\widehat{S}$ be any realization of $\mathbf{S}$ that satisfies the uniform convergence property. Let $K$ be the number of times sets $I_{\mathrm{current}}$ Algorithm 1 (for $\mu$, $\mathsf{minwidth}$, $R(t)=\frac{\alpha T}{t}$) on $\widehat{S}$. Let $t_{i}$ for $1\leq i\leq K$ be the days when $I_{\mathrm{current}}$ is set, where $t_{1}=1$, and we define $t_{K+1}=T+1$ for convenience. We refer to the window $t_{i},t_{i-1})$ as phase $i$.
Let $M^{\widehat{S}}$ be the number of mistakes that Algorithm 1 makes on $\widehat{S}$. Let $M^{\widehat{S}}_{i}$ be the miscoverage rate of Algorithm 1 over phase $i$ of $\widehat{S}$. Let $I_{i}$ be the value of $I_{\mathrm{current}}$ on phase $i$. If Algorithm 1 resets $I_{\mathrm{current}}$ on day $t_{i+1}$, it means that $I_{i}$ has error rate $\leq\frac{\alpha T}{t_{i+1}-1}$ over the window $1,t_{i+1}-1]$ (Algorithm 1, Line 3). By the uniform convergence property (9),